OpenStudy (anonymous):

Can anyone help me understand how to solve this problem? -2x+6i=(-24y)I-14

OpenStudy (anonymous):

By the way, it's real numbers and imaginary numbers. I know the steps except the "i" next to (-24y) throws me off

OpenStudy (anonymous):

-2x+6i=(-24y)I-14 you distribute the i -2x+6i=-24yi-14 -2x=-24yi-14-6i subtract 6i from both sides x=12yi+3i+7 divide everything by -2

OpenStudy (anonymous):

@hector1 the answer on the text book is 1 and -1 but I have no idea how :( lol

OpenStudy (anonymous):

Oh sorry! I was looking at the wrong one. But thank you